OpenStudy (anonymous):

Which number is irrational? A. square root of 64 B.square root of 256 C. square root of 100 D. square root of 80

OpenStudy (anonymous):

@jabezz177

OpenStudy (anonymous):

irrational means that it can't be expressed as a fraction, and if you find the square root of each then you will see that one of them cannot be expressed as a fraction square root of 62 = 8 = 8/1 sqrt(256)=16=16/1 sqrt(100)=10=10/1 sqrt(80) which is about 8.94427 can't be written as a fraction, so the answer is D. 80

OpenStudy (anonymous):

cool thx